### v1

# Computational Sentence-level Metrics Predicting Human Sentence Comprehension

Abstract
The majority of research in computational psycholinguistics has concentrated on the processing of words. This study introduces innovative methods for computing sentence-level metrics using multilingual large language models. The metrics developed—sentence surprisal and sentence relevance— are tested and compared to validate whether they can predict how humans comprehend sentences as a whole across languages. These metrics offer significant interpretability and achieve high accuracy in predicting human sentence reading speeds. Our results indicate that these computational sentence-level metrics are exceptionally effective at predicting and elucidating the processing difficulties encountered by readers in comprehending sentences as a whole across a variety of languages. Their impressive performance and generalization capabilities provide a promising avenue for future research in integrating LLMs and cognitive science.
1 Introduction
In the study of how humans comprehend and process language, computational models play a crucial role in understanding the connections between linguistic features and behavior/neural signals. These models can be used to make linguistic predictions, model language features, or specify the processing steps that can be quantitatively compared to behavioral and neural signals.With the abundance of experimental data available, researchers have developed several computational models/metrics to simulate how humans comprehend a given linguistic unit within a given context.
One such model measures the information communicated by any particular linguistic component, whether it is a phoneme, a word, or an entire utterance, when considered within its left context. This method is commonly known as surprisal Crocker et al. (2016). Word surprisal estimates the information among words in context, and this metric has proven to be influential in predicting human word processing (Hale, 2001; Levy, 2008). In contrast to expectation-based models, memory-based models rely on a memory mechanism to store and retrieve information from previous input (e.g., ACT-R (R Anderson, 1975, short-term memory (Baddeley, 2010)). One metric used to estimate language processing involves assessing the semantic relation with other words, such as through semantic similarity (Mitchell and Lapata, 2010; Hollis and Westbury, 2016).
Word surprisal, an expectation-based metric, provides empirical supports for the position that words are more difficult to process when they are harder to anticipate from preceding context (Demberg and Keller, 2008; Smith and Levy, 2013; Hale, 2016; Shain et al., 2020). However, recent work has shown that surprisal computed by neural language models tends to underpredict human reading times of both targeted constructions and naturalistic text (Van Schijndel and Linzen, 2021; Arehalli et al., 2022; Oh and Schuler, 2023). Does this imply that surprisal has lost its efficacy in predicting reading time? Perhaps it is time to reevaluate this conventional metric. On the other hand, the memory-based theory posits processing difficulty arises from storing, retrieving, and integrating previous context with new input. For example, while the dependency-locality model explains certain syntactic structures processing (Gibson, 1998; Vasishth and Lewis, 2006), it has not been as effective as surprisal in predicting humans’ contextual word processing. Another memory-based metric, semantic similarity, gauges the similarity between the meanings of two words or phrases and is effective in predicting how words are processed. Contextual semantic similarity, which was developed from semantic similarity, concerns the semantic relatedness between a target word and its contextual words. Substantial empirical evidence supports the effects of word-level semantic similarity in language comprehension (Roland et al., 2012; Broderick et al., 2018; Sun et al., 2023a; Sun, 2023).
Although the related studies have claimed that some given computational metrics are predictive of sentence processing by humans, the fact is that these metrics can only predict how words are processed in a given context (such as word surprisal) (Hale, 2016; De Varda and Marelli, 2023; Gotlieb Wilcox et al., 2023). However, focusing solely on word processing may not truly mirror the real-world human experience of understanding sentences. In other words, we should focus on how humans comprehend entire sentences, rather than merely the individual words within them. It is necessary to broaden our understanding of sentence comprehension, placing human sentence processing within a larger context: discourse. Sentence-level metrics (such as sentence probability, sentence semantic relevance) are potentially useful in understanding entire sentence comprehension and processing by humans. For instance, the metric of contextual sentence-level relevance computed by Transformer-based language models could predict how Chinese natives comprehend sentences as a whole (Sun and Wang, 2022). A similar method can be applied in investigating sentence comprehension in other languages. In other words, we can enhance our analysis from word-level metrics to sentence-level metrics, exploring their ability to predict the processing of a sentence within a discourse-level context.
Further, during naturalistic discourse reading, a number of factors may have an effect on sentence processing, including expectations concerning the next sentence based on the preceding context, and memory integrated with new input based on semantic relatedness with the context. Surprisal reflects the amount of information conveyed by a linguistic unit and its predictability, while semantic relevance represents the relatedness of a linguistic unit to the other units. Both factors could independently influence processing difficulty, but they may also interact, creating complex processing situations. The study aims to test hypotheses about the nature of these factors and understand their potentials in human sentence comprehension.
With the advent of deep learning techniques and the availability of massive datasets, there is now an opportunity to develop and compute such sentence-level metrics. By using multi-lingual large language models (LLMs), we can estimate the probabilities of the next sentence (sentence surprisal) or compute semantic relevance among sentences (sentence relevance). These sentence-level metrics we proposed are expected to predict how humans comprehend and process sentence as a whole. Because of sentence length and human memory-stored limit, we consider a limited number of surrounding sentences as the context in calculating contextual information and further incorporate the contextual information in our metrics. The method is “attention-aware” approach because it is inspired by the attention mechanism in Transformer (Vaswani et al., 2017).
We plan to apply these computational metrics to test their predictability in a number of languages rather than in English. Fortunately, such datasets and computational tools are available. For instance, existing multi-lingual LLMs are highly capable of processing and understanding text in multiple languages, further helping compute sentence-level metrics we propose. There are also multilingual databases on language comprehension and processing. For instance, the Multilingual Eye-movement Corpus (MECO) is a collection of eye-tracking data that has been collected from participants reading texts in 13 languages (Siegelman et al., 2022). This is an ideal testing dataset for our research. The present study aims to predict sentence comprehension by using sentence-level computational metrics. By employing these metrics, we expect to obtain more accurate predictions of how sentences are processed and comprehended holistically. To evaluate the generalizability of our approach to various languages, we will undertake a multi-lingual investigation.
2 Related Work
Word surprisal, the negative logarithm of a word’s probability in left context, is effectively used in testing expectation-based theories and predicting language processing (Hale, 2016). This metric signifies that predictable words require less cognitive effort to process. However, it remains an open question: which context aspects - lexical, syntactic, semantic, or conceptual - facilitate prediction or integration. As the language models developed so quickly, the methods for calculating word probabilities have been revolutionized in the transition from n-gram to Transformer (Wolf et al., 2020). For instance, recently deep-learning models have dominated many NLP tasks and LLMs have been used as models for calculating the processing difficulty (Goodkind and Bicknell, 2018; Wilcox et al., 2020; Schrimpf et al., 2021). The recent LLMs are effective and convince to compute the probabilities of the next sentence in context. Moreover, “surprisal” is characterized as the negative logarithm of the probability of an event, and surprisal can be applied across different linguistic levels, including sentences, discourse, or phonemes (Venhuizen et al., 2019; Pimentel et al., 2021; Gwilliams and Davis, 2022). Thus, provided we have sufficient computational resources, we can accurately calculate sentence-level surprisals for mutiple languages.
Additionally, some effective methods have been proposed for computing semantic relevance among words within a sentence. The new approach involves integrating contextual information using the attention-aware method, resulting in more powerful contextual semantic relevance (Sun et al., 2023a; Sun et al., 2023b). However, we believe that the new method can be extended and modified to process entire sentences rather than merely words. By employing an attention-aware approach, we can calculate the semantic relevance among sentences, enabling us to measure the extent to which a target sentence is semantically connected to its neighboring sentences within the discourse.
Reading difficulties stem from sources such as word-level factors, syntactic structure, and sentence-level elements. While context significantly affects sentence processing, effective computational metrics are needed to evaluate this. Sentence comprehension involves the mental processes when understanding language utterances, significantly influenced by context and other factors. Although studies have examined various aspects of comprehension such as word processing and syntactic integration (Carpenter et al., 1995; Kamide, 2008; Altmann and Mirković, 2009 Rohde, 2002; Hale, 2016; Gibson, 1998; McRae and Matsuki, 2013), there is a dearth of computational methods to measure context impact and research on comprehending sentences as a whole. This study aims to address this gap by developing computational sentence-leve metrics for evaluating entire sentences across languages.
Specifically, entire sentence comprehension and processing may be influenced by factors like sentence expectations from previous context and memory integration based on semantic relatedness. Word-level surprisal and semantic relevance have been shown to independently affect processing words in context. Words that are both semantically relevant and surprising may create more complex processing situations, requiring more cognitive resources. This study aims to explore whether sentence-level surprisal and semantic relevance are interactive in processing sentences as a whole. Moreover, sesearch on computational approaches to cognitive sciences has predominantly focused on the English language, leading to a lack of investigation in other languages to test the generalizability of findings on surprisal prediction effectiveness (Blasi et al., 2022). It is essential to test hypotheses in multiple languages to enable cross-lingual generalizations using computational models or metrics.
3 Materials and Methods
3.1 Testing datasets
Multilingual databases facilitate cost-effective and reliable cognitive and neural language testing. The Multilingual Eye-tracking Corpus (MECO) is particularly useful due to its diverse range of 13 languages and eye-tracking data indicating cognitive effort in processing (Siegelman et al., 2022). Each language’s participants read 12 encyclopedia entries of 2000 tokens, resulting in 36000 tokens in total. These entries were similarly complex across languages and generated about 70,000 to 80,000 eye-tracking data points per language. MECO was chosen for its language diversity, large native speaker counts, and ample eye-tracking data. Reading speed (or rate) has been extensively explored, typically employing text or individual sentence as the unit of interest (Miller and Coleman, 1971; Carver, 1976; Biancarosa, 2005; Brysbaert, 2019; Siegelman et al., 2022). This study focuses on sentence reading speed (= the number of words in this sentence / total fixation duration of a sentence) - word number per second. When the reading speed is low, readers use more time to process this sentence. In contrast, a higher sentence reading speed indicates that less time is used to process every word for this sentence. While reading speed is not a direct oculomotor measure, it is influenced by oculomotor factors such as fixation duration, saccade durations, and the tendency to skip words. Reading speed is interconnected with these oculomotor measures, as they collectively contribute to how quickly and efficiently one can read and understand a sentence or text. In this sense, reading speed is a broader measure of language comprehension and processing. The reading speed is therefore considered as the overall difficulty or speed when readers comprehend a sentence/text as a whole.
3.2 Computing sentence surprisal
We employed two multilingual LLMs: m-BERT and mGPT to compute sentence surprsial. Sentence surprisal is the negative logarithm of next sentence probability ($-log(p(sentence|left\quad context))$), which is similar to word surprisal. Multi-lingual BERT (m-BERT) (Devlin et al., 2018) can be employed to compute word-level or sentence-level surprisal for various languages. BERT, a masked language model, has good performance in a number of NLP tasks (Salazar et al., 2019; Kalyan et al., 2021), and can be used to estimate next word probabilities. We used the state-of-the-art BERT model (i.e., multilingual-bert-uncased) because it can be consistently applied in different languages in order to compute surprisal. However, it seems a little difficult to use it to compute sentence probability immediately using m-BERT. Fortunately, we can take some strategies to allow m-BERT to approximate next sentence probabilities. For instance, we gauged the joint probability of an entire sentence conditioned on its preceding context. This is achieved by using the chain rule of probability, breaking the sentence into individual tokens, and computing the probability of each token given all preceding tokens (and the context).
Moreover, GPT is essentially an autoregressive model based on the Transformer architecture, trained on a language modeling task, where the objective is to predict the next word in a sequence given the preceding words. The model learns to assign probabilities to words based on the context. The method of chain rule can be introduced in applying mGPT to compute sentence surprisal Shliazhko et al. (2022). The only difference is that GPT employs the language modeling to compute word probability, but BERT uses masked language model to calculate word probability.
To compute sentence surprisal, we also proposed the other methods to compute sentence surprisal. The three methods are summarized as follows: the first one is to use chain rule (CR), which was mentioned above. Sentence surprisal is calculated by tokenizing a sentence into individual elements and computing the joint probability of the sentence given a context. This is done by multiplying the conditional probabilities of each token given its preceding tokens and the context, applying the chain rule of probability to language sequences. The method can be computed based on either m-BERT or mGPT. The second method is to use “next sentence prediction” (NSP mechanism in m-BERT to compute sentence surprsial. Third, sentence surprisal is also quantified by computing the Negative Log-Likelihood (NLL) of the sentence when conditioned on its context. NLL is used as a loss function to measure how well the model’s predictions align with the actual data. For a given sequence of tokens, NLL is a measure of how surprised the model is by the actual sequence. The details on the three methods are seen in Appendix A.
Overall, we employed three distinct methods for computing sentence surprisal, utilizing two multilingual LLMs: m-BERT and mGPT. “ Sentence surprisal” quantifies the level of unpredictability or information associated with encountering a given sentence. The CR and the NLL approach were applied to both LLMs for this purpose. In contrast, the NSP method was exclusively implemented on m-BERT. The methods and models are shown in the Panel A of Fig. 1.
3.3 Computing attention-aware sentence relevance
Figure 1: The computational methods in the present study
This section elaborates on the computation of sentence semantic relevance. Consider a window comprising four sentences, as depicted in Panel B of Fig. 1, where the objective is to evaluate the semantic relationship of the target sentence with the adjacent three sentences. The computation unfolds in two primary steps. Initially, embeddings for each sentence within the window are generated using either m-BERT or mGPT. Subsequently, the application of cosine similarity to the sentence embeddings derived from m-BERT or mGPT serves as a conventional approach to ascertain the semantic similarity between any pair of sentences, as detailed in Appendix B. Nonetheless, this step is to obtain the relevance of two sentences in this window. Our objective, however, is to compute the semantic relatedness between a target sentence and its contextual sentences, which is another sentence-level metric.
Upon calculating the semantic similarity values for any pair of sentences within this window, it is necessary to employ an “attention-aware” approach to process these values across all sentences. The “attention-aware” approach has been successfully applied in computing word-level metrics for predicting reading (Sun et al., 2023a; Sun et al., 2023b). Specifically, we look at the preceding two sentences and the following sentence. The similarity score for any two sentences is computed by taking the dot product between the first embedding of the concatenated input and the second embedding of the concatenated input. Subsequently, we derived four cosine similarity values for these sentences. Each value is then weighted differently based on its distance from the target sentence. It is formalized as (1).
$atten\_sentrev=\text{similarity}_{(s,c)}\cdot W_{(s,c)}\qquad(1)$
Here “W” represents weights according to the positional distance between the target sentence (“s”) and its contextual sentences (“c”). The range of weights is between 1/2 and 1/3 (i.e the neighboring sentence is multiplied by 1/2, but the non-neighboring one is by 1/3.). As the distance from the target sentence is large, the weight is given with smaller one. The weights used for attention-aware metrics mimic human forgetting mechanism, a representation of memory retention decline over time, where retained information halves after each day within a span of several days (Loftus, 1985), shown in Fig. 3. Unlike the attention weights in transformers, which are computed via neural networks and are not easily interpretable. Each cosine similarity value is multiplied by its corresponding weight, and the results are then aggregated to produce a single value. This value signifies the semantic relevance of the target sentence within its context(i.e., “sentence relevance”). Leveraging various sentence embeddings created by m-BERT and mGPT, we are able to derive two distinct metrics of “sentence relevance” for the identical target sentence within its specific context. The computation is shown as Panel B of Fig. 1.
The method proposed for computing attention-aware metrics works as a memory agent (see Appendix B). This approach considers both preceding and following sentences as potential sources of contextual information. The relationship between these contextual sentences and the target sentence is then weighted based on their positional distance and aggregated to create a comprehensive measure of semantic relevance within the discourse. This approach is highly explainable from linguistic and cognitive perspectives. Table 1 provides an overview of the metrics in our analysis.
Table 1: The models and statistical analysis used in this study
Method Measure Equation Statistical Analysis
sentence surprisal chain rule (CR), next sentence prediction (NSP), negative loglikelihood (NLL) $-\log_{2}p(\text{sentence}\,|\,\text{left context})$ GAMM
attention-aware sentence semantic relevance $\sum{C_{(s,c)}\cdot W_{(s,c)}}$ GAMM
3.4 Statistical method and model comparison
We employed Generalized Additive Mixed Models (GAMMs) (Wood, 2017) to investigate the predictive power of certain measures on eye-movements during reading. The performance of GAMM models was evaluated using difference in AIC (Akaike’s Information Criterion) (i.e.,$\Delta$AIC ) between a base GAMM and a full GAMM served as a measure to assess the effectiveness of a metric. A smaller $\Delta$AIC value indicates that the measure we are interested in estimates provide more accurate predictions of the response variable compared to the baseline model.The difference in AIC between a base GAMM and a full GAMM served as a measure to assess the effectiveness of a computational measure. A smaller (negative) $\Delta$AIC
value indicates that the measure we are interested in estimates provide more accurate predictions of fixation duration compared to the baseline model.
Comparing GAMM models, a larger model will have a higher likelihood due to more parameters. But, comparing loglikelihoods (or log-likelihood (LogLik)) directly is not suitable for differently sized models. Both AIC and Bayesian Information Criterion (BIC) address this by penalizing model complexity. However, AIC is preferred for prediction quality, while BIC prioritizes parsimony, often favoring smaller models. Previous studies (Wilcox et al. (2020), Oh and Schuler (2023)) used $\Delta$LogLik to analyze the predictive power of surprisal computed various language models. However, LogLik may not be a good standard. Moreover, including random variables (e.g., participants) and significant predictors like “word length” and “word frequency” is crucial for accurate regression model evaluation. Failure to do so may result in suboptimal fittings for studying reading time. Without optimization, there is a risk of encountering misleading issues that can undermine the validity of the results. The regression models in these studies (Wilcox et al., 2020; Oh and Schuler (2023)) may not be optimal because of lack of random variables or control predictors. The present study included these control predictors (word length/frequency) and random variable (e.g.,“languages”, “participants” ) in the GAMMs. Furthermore, the excessive imposition of heavy penalties can lead to overfitting in GAMM models. We have avoided this practice. More details are seen in Appendix C.
4 Results
4.1 Overall performance
We fitted eight GAMM models to analyze these sentence-level metrics as the main predictors of the response variable (sentence reading speed). The GAMM models also include mean word length and mean word frequency 111“mean word length” refers to the average length of all words within a sentence (i.e., the sum of length of all words / word number), while “mean word frequency” represents the average of the sum of normalized frequencies for all words within a sentence (i.e., the sum of normalized word frequencies / word number). The word frequency information across languages is obtained from https://opus.nlpl.eu/index.php.(Lison and Tiedemann, 2016), and participant as a random effect. This is what an optimal GAMM formula looks like: reading_speed $\sim$ s (mean_word_length), s(mean_log_wordfreq) + s(computational metric, language, by = languages) + languages + s (participant, bs=“re”), data=data (here, s = smooth; ‘by = languages’, it facilitates the inclusion of responses at each language level in the model as distinct terms. Concurrently, smooths by language are employed to capture and model the nuanced variations around these language-specific responses; re = random effect, random slope adjusting the slope of the trend of a numeric predictor). The base GAMM fitting is depicted as: reading_speed $\sim$ s (mean_word_length), s(mean_log_wordfreq) + languages + s(participant, bs=“re”). The $\Delta$AIC represents the difference in AIC values between a full model and a base model.
First, we examine whether a variable is significant or not. When a variable in GAMM is significant, its p-value is smaller than 0.01. These GAMM fittings show that the control predictors, namely mean word length and mean word frequency, are consistently significant across all cases. This indicates that the comprehension of sentences in the context of naturalistic discourse reading is significantly shaped by both the average length and frequency of the words employed. Specifically, the mean word length exerts a negative influence on the speed of sentence reading; in essence, sentences composed of longer words are read more slowly. Conversely, the mean word frequency positively impacts reading speed, meaning that sentences containing words that are more frequently used facilitate faster reading. Therefore, when the average frequency of words in a sentence is high, it enables readers to process the text more swiftly. The effects of word length and frequency are particularly notable in word processing during naturalistic discourse reading. As such, sentence processing and word processing share a great deal of similarity. GAMM fitting results also show that the majority of the metrics we proposed to compute by m-BERT or mGPT are capable of predicting the reading speed data quite well. The random effect of the participant is significant in all GAMM models.
We further compared the performance of the GAMM fittings with different metrics. Table 2 presents the results on $\Delta$AIC for comparing these GAMM fittings. Lower $\Delta$AIC value indicates a better GAMM fitting. It also suggests that the computational metric performs better other. The basis for comparison is the consistent data point numbers (n = 34571) and identical elements in each GAMM fitting. As illustrated in Table 2, the result reveals that sentence surprisal, as calculated using three distinct methodologies — namely, CR, NLL, and NSP — proves to be effective in predicting reading speed. Among the evaluated metrics of sentence surprisal, the performance of m-BERT, when applying the chain rule, stands out as the most effective. Conversely, sentence surprisal calculations using mGPT with NLL were found to be ineffective. The computation of sentence relevance, whether through m-BERT or mGPT, demonstrates viability in prediction accuracy. Notably, the integration of both sentence surprisal and sentence relevance into the GAMM significantly enhances predictive performance, surpassing the results achieved when these metrics are applied independently.
Table 2: $\Delta$AIC for GAMM fittings with different computational measures on the MECO (n = 34571) (CR = Chain Rule; NLL = Negative Log-Likelihood; NSP = Next Sentence Prediction)
GAMM fittings (sentence reading speed) $\Delta$AIC
sentence surprisal (m-BERT-CR) -670.06
sentence surprisal (m-BERT-NSP) -119.94
sentence surprisal (m-BERT-NLL) -506.85
sentence surprisal (mGPT-CR) -393.56
sentence surprisal (mGPT-NLL) NA
attention-aware sentence relevance (m-BERT) -445.86
attention-aware sentence relevance (mGPT) -294.1
sentence surprisal (m-BERT-CR), sentence relevance (m-BERT) -1397.6
Figure 2: The overall partial effects of the primary predictors—sentence surprisal and sentence relevance—on reading speed across languages. The x-axis denotes the metric, and the y-axis represents the reading speed. Sentence surprisal and sentence relevance are transformed by logarithm in order to be get closer normal distribution, further having better fittings. Each curve depicts the relationship between a predictor variable and the response variable, reading speed. Steeper slopes on these curves indicate a stronger relationship between the predictor and reading speed, while flatter slopes suggest a weaker effect.
The overall partial effects of the metrics of our interest (excluding the “languages” factor) in GAMM fittings are shown in Fig. 2. The results reveal that sentence surprisal (by m-BERT and CR basically adversely affects reading speed across 13 languages, indicating that a higher level of surprisal for a sentence is associated with slower/lower reading speed. On the other hand, sentence relevance (by m-BERT) positively influences reading speed across these languages, suggesting that greater relevance for a sentence facilitates higher/faster reading speed.
4.2 Performance in individual languages
The results from fitting the GAMM offer valuable insights into the importance of the variables considered. Using the $\Delta$AIC, we can determine the relative performance of different GAMM configurations. We identified the optimal metrics for our study: sentence surprisal calculated using m-BERT with the chain rule (CR), and sentence relevance determined by m-BERT. This analysis enables us to select optimal GAMM structures for evaluating how computational metrics predict sentence reading speed across 13 languages. The optimal GAMM for each language is specified as follows:reading_speed $\sim$ s (mean_word_length), s(mean_log_wordfreq) + s(sentence  surprisal) + s(sentence  relevance) + s (participant, bs=“re”), data= language_data. To calculate the $\Delta$AIC for “sentence relevance”, we excluded $s(sentence\quad surprisal)$ from the model, comparing the AIC of the complete model with that of the model absent “sentence surprisal”. Conversely, to determine the $\Delta$AIC for sentence surprisal,$s(sentence\quad relevance)$ is removed, and the AIC of the full model is subtracted from the model without “sentence relevance”. The significance of either “sentence surprisal” or “sentence relevance” is assessed using the p-value ( at p-value of 0.01) and the shape of the curve. These metrics were selected to represent, respectively, the concepts of sentence surprisal and sentence relevance in our examination of each language. The results are illustrated in Fig 4 in Appendix C. It appears that sentence surprisal lacks significance in English, Korean, and Russian, while sentence relevance shows no significant impact in English, Russian, Spanish, and Turkish.
We employed T-test to assess the statistical significance of the $\Delta$AIC values for both sentence surprisal and sentence relevance. The analyses revealed no significant differences, with the p-value substantially exceeding 0.05. This outcome suggests that sentence surprisal and sentence relevance may contribute equivalently to the prediction of reading speeds. Consequently, it can be inferred that both metrics are effective in facilitating the overall comprehension of sentences. Despite this, Table 2 shows that the overall predictive power of sentence surprisal with m-BERT and CR is stronger than sentence relevance. Moreover, sentence surprisal and sentence relevance are totally distinct metrics, and their overall correlation is -0.054 (p-value being zero), their correlation in each language is also remarkably small, seen in Appendix D.
5 Discussion
Word surprisal has been found to predict language comprehension across various datasets (Ryskin and Nieuwland, 2023). Sentence prediction has also been found in impacting language comprehension(Goldstein et al., 2022; Kriegeskorte et al., 2008). However, sentence-level computational metrics have not been proposed previously. Our study shows that the surprisal of a sentence can predict and clarify the difficulties readers encounter in comprehending a sentence as a whole. If a sentence is highly predictable given its left context, readers may feel easier to comprehend or process, leading to higher/faster reading speed. On the other hand, if a sentence is more unpredictable (i.e., lower sentence surprisal), readers may be more likely to experience processing difficulty. This can lead to slower/lower reading speed and potentially disrupt the flow of reading. The ability to predict the next sentence may be a fundamental aspect of language comprehension, as it allows users to anticipate upcoming information, and to integrate it smoothly with the preceding information. The impact of sentence surprisal on reading aligns closely with the influence of word surprisal on the reading process(Goodkind and Bicknell, 2018; Gotlieb Wilcox et al., 2023). Prediction occurs not only at the level of individual words, but also at the level of entire sentences in discourse. Prediction could be a key mechanism underlying human language comprehension, which involves using the context and prior knowledge to make predictions about the likely content of the upcoming input, and adjusting those predictions as new information is encountered.
Numerous studies have also identified the influence of context on sentence processing (Cohen and Servan-Schreiber, 1992; Grisoni et al., 2017). However, we are the first to propose to compute sentence-level semantic relevance. The sentence relevance computed by attention-aware method can also effectively help interpret and predict the processing difficulties readers face when comprehending sentences in their entirety. When a sentence is more semantically relevant to the context (in discourse) in which it appears, it is more likely to be processed quickly and accurately, and readers are more likely to understand its meaning without having to invest a lot of cognitive effort (i.e. higher reading speed). Conversely, when a sentence is irrelevant with the context (i.e., lower sentence relevance), it may slow down reading speed and require more cognitive effort. The impact of sentence relevance on reading closely resembles the effect of word semantic relevance on the reading (Sun et al., 2023b; Sun et al., 2023a). There are several reasons why relevant sentences tend to be processed more quickly. According to the research of discourse structure, relevant sentences can easily create discourse coherence. One key factor is the activation of mental representations related to the topic or theme of the text (Traxler, 2011). When a sentence is relevant to the context, it could activate mental representations through memorizing the context, making it easier for readers to process the sentence. This activation can lead to faster reading speed, as readers are able to effectively activate the memory of the context and quickly integrate it into their mental representation of the text. In contrast, irrelevant sentences may require more cognitive effort to process, as they could not activate the memory from the context and do not fit smoothly into the mental representation of the discourse that readers are constructing. This processing difficulty can slow down reading speed.
The expectation and memory could be mutually interacted (Ryskin and Nieuwland, 2023) in language comprehension/processing. Sentence surprisal measures the processing difficulty of a sentence by assessing its unexpectedness relative to the preceding context. On the other hand, sentence relevance represents memory by gauging how well a sentence fits into the context, with contextually relevant sentences being easier to process, resulting in faster reading speed. By combining them, it can encompass both processing difficulty, and could better simulate human language processing.
6 Conclusion
This study presented sentence-level metrics for predicting human comprehension of sentences as a whole. The results show that both sentence surprisal and sentence relevance were highly capable of predicting human sentence prediction. The methods of computing sentence surprisal worked well. The attention-aware method allowed for computing contextual information for sentence-level semantic relevance. All of these methods also exhibited strong generalization capabilities across languages. The findings showed that these sentence-level metrics are informative features for modeling human sentence comprehension across languages. Our work highlighted the potential of combining computational models with cognitive models to better explain and predict human language comprehension/processing, and further to develop more effective NLP and AGI systems.
7 Limitations
The study introduces an innovative approach to predicting sentence comprehension using attention-aware computational metrics. However, it comes with several limitations. Primarily, the methodology depends on specific multilingual LLM (m-BERT and mGPT). The variation in proficiency across different languages can result in disparate abilities to process these languages. In essence, sentence surprisal or sentence relevance might not be estimated with equal precision across all languages using multilingual LLMs. This limitation could affect the applicability of our metrics across varied linguistic structures and cultural contexts, leading to varying performance of our metrics in different languages. Moreover, the attention-aware method, while effectively capturing some aspects of contextual information, may oversimplify the complex dynamics of human cognitive mechanisms on language comprehension and processing. Addressing these limitations could pave the way for more comprehensive models that better mirror the intricacies of human language comprehension and processing.
Data Availability
A. Computing sentence surprisal
This section details how to compute sentence-level surprisal using the three methods: chain rule (CR), negative log-likelihood (NLL) and next sentence prediction (NSP). The first two methods were applied in either m-BERT or mGPT, and the third method was merely implemented in m-BERT.
The following provides a detailed account of how to apply chain rule (CR) to compute sentence surprisal . First, we tokenized sentence into various tokens: $S=[t_{1},t_{2},\ldots,t_{n}]$. Probability of a token $t_{i}$ in the sentence S given its left context C in a text and preceding tokens can be represented as: $P(t_{i}|C,t_{1},t_{2},\ldots,t_{i-1})$. (S = sentence, t = token, C = context). Based on this, we can calculate the probability of each token prior to its left context. After obtaining the probability of each word, we computed the joint probability of the entire sentence in equation (2): S given its context C, $P(S|C)=P(t_{1}|C)\cdot P(t_{2}|C,t_{1})\cdot P(t_{3}|C,t_{1},t_{2})\cdot\ldots\cdot P(t_{n}|C,t_{1},t_{2},\ldots,t_{n-1})\quad(2)$, further getting sentence surprisal, $-log(P(S|C))$. The equation is an application of the chain rule of probability to sequences (like sentences). It captures the idea that the probability of a sentence given a context can be decomposed into the product of conditional probabilities of its individual tokens. The following provides the details why it works. When you have a sequence of events or tokens, the chain rule can be extended. For a sequence of three events $t_{1}$, $t_{2}$, and $t_{3}$: $P(t_{1},t_{2},t_{3})=P(t_{1})\times P(t_{2}|t_{1})\times P(t_{3}|t_{1},t_{2})$. In the case of natural language, sentences often have a context. The context can be prior knowledge, a preceding sentence, or any other relevant information. Given a context (C), the probability of a sequence changes (S): $P(S|C)=P(t_{1}|C)\times P(t_{2}|C,t_{1})\times P(t_{3}|C,t_{1},t_{2})\times\ldots\times P(t_{n}|C,t_{1},t_{2},...,t_{n-1})$. The method of computing sentence probability is consistent with the principles of probability theory and the way sequence modeling is approached in the context of NLP.
Moreover, using Negative Log-Likelihood (NLL), we can also compute the surprisal of a sentence conditioned on its preceding textual content. The specific procedure is detailed as follows: the text is tokenized into a sequence of token IDs, converted into a PyTorch tensor, and processed through the model to determine the NLL. NLL is used as a loss function to measure how well the model’s predictions align with the actual data. For a given sequence of tokens, NLL is a measure of how surprised the model is by the actual sequence. There are two steps in calculation: 1) Sum up these NLLs for the words in the sentence: $\text{NLL}(S\big{|}C)=−\sum\log(P(t_{i}|t_{1},t_{2},...,t_{i-1})$ ($t_{i}$ is the word in the target sentence “S”, and $t_{1}$ , $t_{2}$ , …, $t_{i-1}$ include both the context and the preceding words in the sentence) 2) Convert NLL to Probability, shown in equation (3): $P(S|C)=e^{-\text{NLL}(S|C)}\quad(3)$. The term $e^{-\text{NLL}}$ represents an exponential function with the negative Natural Log Loss (NLL) of $S$ given $C$ as the exponent.
In addition to adopting the chain rule, we could employ the “next sentence prediction (NSP)” mechanism in BERT to compute sentence surprisal. To compute the probability that one sentence follows another using BERT’s NSP, begin by preparing the input. Combine the two sentences, placing a special [CLS] token at the start and a [SEP] token in between them. After tokenizing this combined sequence with BERT’s tokenizer, pass it through the BERT model. BERT uses the representation of the [CLS] token, which encapsulates information about the entire input, to predict the relationship between the two sentences. The model outputs two probabilities: “IsNext” and “NotNext”. The “IsNext” probability indicates the likelihood that the second sentence logically follows the first. By examining this probability, one can gauge how likely the model perceives the given sentence order to be. The above description can be summarized as the following two equations. Probability of sentence B being the next sentence given the left context A is formalized as (4):
$P(B|A)=\text{softmax}(w_{2}^{T}[\text{CLS}]+b_{2})\qquad(4)$
where [CLS] denotes the representation of the first sentence obtained from the final layer of the BERT model, and the softmax function computes the probability distribution over all possible next sentences. $w_{2}$ and $b_{2}$ are learnable weight vector and bias term, respectively. And then we can obtain the sentence surprisal. However, NSP in BERT is a binary classifier, that is, the model determines if a sentence logically follows a given sentence or not. This is a simplification of the rich structure and semantics in natural language. The use of NSP may raise the question about the interpretability of BERT’s estimated probabilities for subsequent sentences, especially when viewed from a cognitive modeling standpoint.
B. Attention-aware approach and its memory capability
This section provides a comprehensive guide on calculating sentence relevance using an attention-aware approach, executed in a two-step strategy. The initial step involves generating sentence embeddings using m-BERT and mGPT, followed by computing the similarity between sentences based on these embeddings. The subsequent step entails applying the attention-aware approach to manage multiple similarity values across several sentences within a window stack.
In the first step, generally, using cosine similarity to compare the sentence embeddings generated by m-BERT or mPGT is a common approach for computing semantic similarity between sentences, formalized as Equation (5).
$\text{similarity}(s,c)=\frac{e_{s}\cdot e_{c}}{\lVert e_{s}\rVert\lVert e_{c}\rVert}\qquad(5)$
[…]
We employed two distinct multilingual LLMs (i.e., m-BERT and mGPT), noting subtle differences in their approaches to generating sentence embeddings. Regarding m-BERT, when a sentence is input into BERT, it is first tokenized and then prepended with a special “[CLS]” token. After processing through BERT’s layers, the embedding corresponding to this “[CLS]” token is often used as the sentence embedding. The tokenized input sentence (S) is represented as [CLS], $t_{1}$, $t_{2}$, …, $t_{n}$. After processing through BERT, the output embeddings at the final layer for this sequence are [$E_{\text{CLS}}$, $E_{t_{1}}$, $E_{t_{2}}$, …, $E_{t_{n}}$]. Mean pooling involves calculating the average of the embeddings of all tokens in the sequence [CLS].
Here is Equation (6) to compute a sentence (or a text) embedding using m-BERT.
$e_{s}=\text{BERT}_{\text{pooler}}(s)=\frac{1}{n}\sum_{i=1}^{n}h_{i}$ (1)
where $s$ is the input sentence, and $\text{BERT}_{\text{pooler}}$ is a mean pooling layer that takes the output vectors $h_{i}$ of all $n$ tokens in the input sentence and computes their average to obtain the sentence representation $e_{s}$.
Moreover, we used mGPT to compute sentence similarity for cross-validation. We still applied embedding-based method to do this. Specifically, mGPT was to obtain embeddings for each sentence. Still employing the mean pooling, we used the hidden states of the sentence to represent the embedding. Cosine similarity was to calculate the similarity between the two embeddings, which is similar to BERT-similarity computation.
The subsequent discussion highlights the distinctions between m-BERT and mGPT in generating sentence-based embeddings. First, mGPT, being primarily focused on generative tasks, does not utilize a special token like [CLS] for aggregating sentence meaning. GPT architecture is designed to predict the next token in a sequence based on the previous context, which inherently focuses on a unidirectional flow of information. The other difference is to use pooling for sentence embeddings. Specifically, to obtain sentence-level embeddings from GPT, one common method is to aggregate the hidden states (from the last layer) of all tokens in the output. Mean or max pooling can be applied to these token-level embeddings to create a single vector representing the entire sentence. This approach leverages the contextual information encoded by GPT in a sequential manner, albeit without the bidirectional context that BERT captures. In BERT, pooling is an alternative to using the “[CLS]” token embedding, offering a way to capture a distributed representation of sentence meaning. However, in GPT, pooling is a necessary step for sentence-level representation since the model lacks a mechanism like the [CLS] token for summarizing the text.
After completing the first step, we elaborated on applying the attention-aware method to calculate sentence relevance. Upon acquiring the similarity values for sentences within the window stack, we proceeded to apply weights to these values. The process involves aggregating the weighted similarity values to derive a final score, as detailed in the main text. Our aim is to underscore the efficacy of the attention-aware approach in capturing contextual nuances and to explain the underlying mechanism of the weighting system.
Fig.3 shows how the attention-aware approach we adopted simulates a short-term memory stack, mirroring how readers retain memory of previously encountered words and their meanings. The method for using various weights of semantic relevance between any two sentences is inspired by both the attention mechanism found in Transformers (Panel A in Fig.3) and the human process of forgetting (Panel B in Fig. 3). The attention-aware approach can facilitate effective incorporation of contextual information and enabling it to achieve memory storage, retrieval, and integration. The weight values gradually decrease with the distance between the target sentence and the contextual sentences (see Equation (1) and the Panel B of Fig. 1), similar to the forgetting curve (Loftus, 1985).
As illustrated in the Panel B of Fig. 1, in a window stack, sentences closer to the target sentence resemble the initial days in the forgetting curve, while more distant words resemble the latter days. To simulate human forgetting mechanism, we allocated larger weights to the closer sentences and smaller weights to the distant sentences and the similarity between human forgetting mechanism and attentional weights adopted in the current study. The attention-aware approach could be linked to memory models in terms of how memory is decayed during the encoding of information, which subsequently affects how humans process sentences during reading. More importantly, the attention-aware approach is computational and fundamentally memory-based, facilitating memory storage, retrieval, and integration. This approach not only realizes the memory function but also incorporates the expectation effect.
Figure 3: The memory capability and weights adopted in the attention-aware approach
C. Statistical methods
To meet our goals of accurately predicting multilingual eye-tracking data, we utilized Generalized Additive Mixed Models (GAMMs)(Wood, 2017). GAMMs are effective in analyzing nonlinear effects and multiplicative interactions between variables, making them ideal for evaluating the predictability of semantic similarity. They are more flexible than traditional regression methods in modeling complex relationships between variables. Eye-tracking data is simpler to analyze statistically than EEG and fMRI data, which makes it an ideal choice for our study on naturalistic discourse reading. However, assessing model performance and comparing models can be challenging, and relying solely on correlations can be limiting. Fortunately, GAMMs are well-suited for comprehensive and precise assessments of model performance. We compared models using AIC (Akaike’s Information Criterion) values, where a smaller value indicates a better model.
AIC or BIC (Bayesian Information Criterion ) are both measures of model fit that balance goodness of fit with model complexity. Lower values of AIC or BIC indicate better model fit. However, AIC is a popular criterion for comparing GAMMs, and it has some advantages over other criteria. AIC is designed to balance the trade-off between model fit and model complexity, penalizing models with more parameters. This makes it useful for selecting models that provide a good balance between fit and complexity. AIC is also relatively easy to compute and widely used in statistical modeling.
Comparing two GAMM (or LMER) models, where one is larger than the other (having all the parameters of the other model and some additional ones), the likelihood will always be higher for the larger model. This is because a larger model can fit the data better by having more parameters. However, directly comparing loglikelihoods is not appropriate when models have different sizes. Both AIC and BIC address this issue by incorporating a penalty for the number of parameters. AIC and BIC have different principles for penalizing complexity. Generally, BIC penalizes complexity more strongly than AIC, tending to favor smaller models unless both approaches agree. In essence, AIC is preferable when the main goal is prediction quality, as a slightly larger model can still provide good predictions, while a too small model usually does not. On the other hand, BIC aims to identify a reasonably sized true model by prioritizing parsimony. BIC is often better in finding the true model, but it has a higher chance of selecting a model that is too small, which is not favorable for prediction. In practice, the true model is often not “small”, but for reasons such as interpretability, smaller models are sometimes preferred even if they have slightly worse prediction performance, in which case BIC may be preferred. In summary, from the perspective of AIC, it is better to fit a slightly larger model than a too small one for improved prediction quality. However, from the perspective of BIC, both excessively large and excessively small models are equally undesirable.
In the studies conducted by Wilcox et al. (2020) and Oh and Schuler (2023), the relationship between model perplexity and $\Delta$LogLik (log-likelihood) was utilized to analyze the perceptual competence of surprisal generated by different LMs. Their objective was to determine which LMs were capable of generating more powerful surprisal based on various corpora. In contrast, the current study aims to assess the predictive performance of our algorithms.
It is important to highlight some potential issues when using $\Delta$LogLik as a criterion. Wilcox et al. (2020) utilized GAMMs for their analysis but did not incorporate any random variables in these models. Consequently, confirming the optimality of these GAMMs becomes challenging. On the other hand, Oh and Schuler (2023) mentioned that they employed Linear Mixed Effects Models (LMER) and included random effects. However, the LMERs used in Oh and Schuler (2023) did not include “word frequency” as a control predictor. It is well-established in psycholinguistics and cognitive science that both “word length” and “word frequency” are significant variables in predicting reading time. Including these two factors as control predictors is commonly practiced when studying reading time. Therefore, when investigating reading time, it is advisable to include both “word length” and “word frequency” in GAMMs or LMERs. Meanwhile, an optimal model should also include random variables. Failure to do so may result in suboptimal GAMMs or LMERs for studying reading time. Moreover, The excessive application of heavy penalties (e.g., k in GAMMs) on the given metrics leads to overfitting in the mixed-effect models. For example, the heavy penalty on the specific metric results in the partial effect curve of this given metric becoming much steeper, and the AIC or likelihood in the model increasing remarkably. Conversely, removing such penalties eliminates these effects.
Considering these factors, the GAMM models used in the present study include two control predictors: the mean word length for a sentence and the mean word length for a sentence. Additionally, random variables, participants in eye-tracking experiments, and languages, were included. We used fs (i.e.,random smooths) to adjust the trend of a numeric predictor in a nonlinear way, which includes random intercept and random slope. In other words, we can allow the metrics of interest to explore their effects on various levels of random variables comprehensively. We employed AIC to compare the performance of different GAMMs. The baseline model excludes the main predictor of interest (sentence surprisal or attention-aware sentence surprisal) but retains the other elements in the full GAMM model. After fitting the regression models, we calculated the $\Delta$AIC
values for each GAMM model by subtracting the AIC of the base GAMM model from that of a full GAMM model. A smaller AIC indicates better model performance. Similarly, a smaller $\Delta$AIC
also indicates better performance. Additioanlly, considering the best sentence surprisal and sentence relevance in each language, we used t tests to check wether $\Delta$AIC
values for two metrics have significant differences. The result shows that there is no significance difference. In other words, it is not easy to distinguish sentence surprisal or sentence relevance could predict reading speed better or not.
Figure 4: The partial effects of the primary predictors—sentence surprisal and sentence relevance—on reading speed across 13 languages (i.e., Dutch, Estonian, Finnish, German, Greek, Hebrew, Italian, Korean, Norwegian, Russian, Spanish, Turkish). Note: The upper section of the diagram features sentence surprisal, while the lower portion is dedicated to sentence relevance.The x-axis signifies the computational metric, while the y-axis delineates the reading speed. To achieve a closer approximation to a normal distribution, and consequently improve the fitting, all metrics undergo a logarithmic transformation. Each curve visually articulates the correlation between a predictor variable and the response variable, namely reading speed. A steeper incline on these curves underscores a more robust impact between the predictor and reading speed, whereas gentler slopes imply a less pronounced effect. Moreover, when a curve fluctuates around zero, its effect vanishes. The information regarding $\rho$-values and $\Delta$AIC is displayed at the top of each plot. The methodology for calculating $\Delta$AIC for “sentence surprisal” and “sentence relevance” is detailed in the main text. In conclusion, sentence surprisal seems to lack significance in English, Korean, and Russian. However, sentence relevance may show no significant impact in English, Russian, Spanish, and Turkish.
D. Correlation between sentence surprisal and sentence relevance
The overall Pearson correlation between sentence surprisal (computed using m-BERT with the chain rule) and sentence relevance (also derived from m-BERT) stands at -0.054. The value suggests that there is a weak correlation between the two metrics, and indicating that the two metrics are completely distinct. This relationship across the 13 languages is depicted in Fig. 5. The observed correlations among the metrics for each language are notably minimal. Such low correlation scores underscore the fact that the sentence surprisal and sentence relevance we calculated represent entirely distinct metrics.
Figure 5: Pearson correlation between sentence-level surprisal (computed by m-BERT and chain rule) and sentence-level semantic relevance (computed based on m-BERT) in each language. Note: the abbreviations for these 13 melange are as follows. du = Dutch; ee = Estonian; en = English; fi = Finnish; ge = German; gr = Greek; he = Hebrew; it = Italian; ko = Korean; no = Norwegian; ru = Russian; sp = Spanish; tr = Turkish